OpenStudy (anonymous):

between which two integers does square root of 12 lie?

geerky42 (geerky42):

HINT: \(\sqrt 9 = 3\) and \(\sqrt{16} = 4\)

geerky42 (geerky42):

maybe too much of a hint...

OpenStudy (anonymous):

i am gonna go out on a limb here and say 3 and 4? :p

geerky42 (geerky42):

Yep, Do you understand why?

OpenStudy (anonymous):

not really...i am horrible at algebra

geerky42 (geerky42):

square root is just inverse of square. ex: \(3^2 = 9\), so \(\sqrt{9} = 3\). \(4^2 = 16\), so \(\sqrt{16} = 4\) So \(\sqrt{12}\) must be in between 3 and 4, right?

OpenStudy (anonymous):

oooook. It makes a little more sense now, i guess. more than it did before. Thanks Geerky